OpenStudy (anonymous):

What is the equation of the line perpendicular to the line y =(4/9) x – 2 that passes through the point (4,3)?

OpenStudy (anonymous):

perpendicular lines have negative reciprocal slopes --> so new line needs slope of -9/4 then use y-y = m(x-x)

OpenStudy (anonymous):

Okay wait... I still need help. Would it be -9/4x+12 or -12?

OpenStudy (anonymous):

y-3 = -(9/4) (x - 4) y-3 = -9/4x + 9 y=-9/4x+12